# Instruction Matters, a Simple yet Effective Task Selection Approach in Instruction Tuning for Specific Tasks

### v1

@@ 5.3 NIV2 Results
Figure 1 illustrates the experimental results for 33 tasks within NIV2666Though NIV2 has 12 task clusters and 119 tasks for evaluation, we randomly select up to 3 tasks per task cluster due to computational cost.. We adhere to [Def + Pos(2)] setting in NIV2, which includes Task Definition and two Positive Task Examples in the instruction, but note that we only use Task Definition for task selection. The precise input format utilized for training and evaluation is elaborated in the Appendix D.3.
@@ 6 Further Analysis
For further analysis, we conduct experiments to compare instruction-based versus data sample-based task selection, examine the impact of instruction refinement in the P3 dataset, and assess the effects of varying the number of related tasks selected (k). Additionally, changes in task selection performance related to the number of instructions and experiments on learning different formats of meta-datasets and their impact on performance are detailed in Appendices I and J.
@@
While the T5(3B) + DSTa model outperforms in specific tasks like Hellaswag, it falls short in others, including RTE and WSC. Conversely, T5(3B) + InsTa consistently enhances performance across the majority of tasks relative to T0-3B, and achieves an impressive average performance enhancement of 5.08% over T5(3B) + DSTa. This outcome suggests that instances within data samples might obstruct the extraction of representative task features, diminishing the task selector’s ability to identify related tasks effectively. Please refer to Appendix G for more experimental details.
@@ 6.2 Impact of Instruction Refinement
As mentioned in Section 4.2, our approach operates based on the premise that the instruction accurately describes the characteristics of task, necessitating the process of instruction refinement. Table 6 demonstrates the impact of such instruction refinement on performance. The term "Unfiltered" denotes the training conducted without instruction refinement, while "Filtered" indicates the use of refined models.
Unfiltered models, while showing performance improvements in most tasks compared to T0-3B, encounter performance degradation in certain tasks, notably StoryCloze, due to the selection of irrelevant tasks. Conversely, models that use filtered and refined instructions accurately select related tasks and generally demonstrate improved performance over unfiltered models. This result emphasizes the significance of instruction quality in task selection. Moreover, the notable performance enhancement seen in T5(3B) + InsTa indicates the substantial role of instruction quality, particularly when additional training for the alignment is held.
@@ 6.3 Scaling Relevant Tasks
We explore how the model performance is affected by the number of selected tasks during instruction tuning. Figure 2 illustrates the average performance of T5(3B) + InsTa and T5(3B) + InsTa, across different numbers of selected tasks (). The performance of the top-35 tasks for P3 and top-756 tasks for NIV2 corresponds to the fully instruction-tuned models, representing the scores of T0-3B and Tk-Instruct-3B, respectively. On average, we observe a progressive improvement in performance up to five tasks for P3 and seventy tasks for NIV2, after which it declines, affirming the adverse impact of non-relevant tasks during training. For more analysis and detailed results, please refer to Appendix H.
@@ 7 Conclusion
In this study, we discover that selecting informative tasks for instruction tuning can be effectively achieved by exclusively using task instructions. Our experiments reveal that this method, particularly when aligned with the meta-dataset’s instructions, surpasses traditional methods which depend on data samples to determine task relevance. Beyond its robust performance, our approach’s most significant advantage is its simplicity: as long as the instruction (task description) of the target task can be described, our method can be applied. This approach marks a significant shift away from traditional methods that require exhaustive pairwise comparisons or the labor-intensive creation of data samples for new tasks. By adopting an instruction-only strategy, our method simplifies and automates the task selection process for instruction tuning, providing a more efficient and practical approach to building models in real-world scenarios.
@@
While highlighting effective task selection method in instruction tuning, we do not perform experimental results over the different sizes of model parameters other than the T5(3B) parameter model due to computational cost. For example, a language model bigger than 11B parameters may be less susceptible to negative transfer due to model capacity, or the effectiveness of our task selection method might stand out even more. Moreover, we only use the encoder-decoder architecture model in the paper, we leave the investigation on the decoder model, such as LLaMA 7B/13B, to our future work.
@@ A.1 Task selection in Intermediate-Task Transfer Learning
Within the scope of intermediate-task transfer learning, the goal of task selection is to enhance the performance by further training the model on related intermediate tasks before fine-tuning it on a target task Phang et al. (2018); Vu et al. (2020b); Kung et al. (2021); Poth et al. (2021b). Therefore, it presupposes the availability of a labeled dataset for the target task..
Vu et al. (2020b) used BERT as a feature extractor to identify related tasks by comparing embeddings with an auxiliary task pool. In contrast, Kung et al. (2021) developed a more efficient method, training a task discriminator with just 500 samples from each dataset, thereby reducing data needs. However, creating 500 labeled samples for a target task can still be burdensome, especially in an instruction-tuning setting. Moreover, in line with Vu et al. (2020b), this approach entails a high computational cost for training task discriminator and inference on all auxiliary instances.
Poth et al. (2021b) opted for a different approach, directly measuring pairwise transfer for all datasets instead of using samples’ embedding. To reduce computational costs, they employ adapter structures. Nonetheless, this method still incurs significant costs in terms of target data construction and computational overhead because it requires training and inference across all models.
@@ A.2 Task selection in instruction tuning
In task selection in instruction tuning, the objective is to select and train related task for improving zero-shot performance for target task. Unlike intermediate-task transfer learning, where samples are composed solely of instances, data samples comprise instructions plus instances. A model trained with such instructions can tackle new tasks given new instructions, utilizing previously learned instructions as a basis.
@@
Ivison et al. (2023) propose a sample-based task selection methodology in instruction tuning. They encode all instances from a vast multitask data pool using a T5-3B and then build a search index from the resulting representations. A key advantage of their approach is that there’s no need to train an embedding generator with the target task’s data samples, unlike task selection in intermediate-task transfer learning. However, there is still a limitation of requiring 1000 unlabeled data samples for the target data. Jang et al. (2023) adopted a method similar to Ivison et al. (2023) to retrieve experts related to the target task for instruction tuning by encoding 32 data samples for each expert to calculate similarity.
@@
Zhou et al. (2023); Kim et al. (2023a) applied a similar methodology of Poth et al. (2021b) to instruction tuning, performing pairwise transfer across a variety of tasks. Notably, Zhou et al. (2023) conducted pairwise transfer for all tasks in the P3 dataset, and these results were utilized as a baseline in their paper.
A.3 Baseline for Task Selection
@@
As noted by Ivison et al. (2023), task selection in intermediate-task transfer learning has several disadvantages compared to the task selection methodologies used in instruction tuning settings. These methodologies are primarily applied to classification tasks, require a large number of labeled samples for the target task, and involve high computational costs for training a model to generate task embeddings. For these reasons, we exclude these studies in our baselines. Additionally, we used Jang et al. (2023)’s methodology as the baseline for the sample-based task selection instead of Ivison et al. (2023) because Jang et al. (2023)’s approach requires fewer data samples, making implementation easier. The characteristics and practical applicability of each methodology are summarized in the Table 7
Appendix B Off-the-shelf Embedding Model
@@
During training, we employ a learning rate of 1e-6 for the P3 dataset and 1e-5 for the NIV2 dataset. The remaining training configurations align with those described in Reimers and Gurevych (2019). Specifically, we conduct training for 5 epochs and select the checkpoint corresponding to the best validation performance for subsequent analysis and evaluation.
Appendix C Training Details
@@
Appendix D Examples and Formulation of Instructions of P3 and NIV2.
D.1 Examples of P3 and NIV2 Instructions.
@@ D.3 NIV2 Input Formats
Appendix E GPT-4 API Generation
E.1 NIV2 Instruction Generation
E.2 Big-Bench Hard Instruction Generation
The BBH tasks don’t have refined Task Definition like NIV2 datasets. In order to select relevant tasks using NIV2 datasets, and to compare the performance with Tk-Instruct, we generate instruction for BBH tasks using GPT-4. Table 13 represents API query used for instruction generation, and Table 14, 15 and 16 show generated 27 instructions for each task.
Appendix F Pairwise Transfer Task Selection vs. Instruction-Based Task Selection
@@ E.2 Big-Bench Hard Instruction Generation
Appendix G Experimental Detail for Data Sample-based Task Selection
@@
Appendix H Analysis on Selecting Top-k Relevant Tasks
@@
Appendix I Scaling The Number of Instructions in Alignment
@@
Appendix J Learning Different Formats of Meta Datasets and Its Impact on Performance

### v2

@@ 5.3 NIV2 Results
Figure 1 illustrates the experimental results for 33 tasks within NIV2666Though NIV2 has 12 task clusters and 119 tasks for evaluation, we randomly select up to 3 tasks per task cluster due to computational cost.. We adhere to [Def + Pos(2)] setting in NIV2, which includes Task Definition and two Positive Task Examples in the instruction, but note that we only use Task Definition for task selection. The precise input format utilized for training and evaluation is elaborated in the Appendix F.3.
@@ 6 Further Analysis
For further analysis, we conduct experiments to compare instruction-based versus data sample-based task selection, examine the effects of instruction refinement in the P3 dataset, and assess the impacts of varying the number of related tasks selected (). Additionally, changes in task selection performance based on the number of instructions used for alignment, as well as experiments on learning different styles of meta-datasets and their impact on performance, are detailed in Appendices K and L.
@@
While the T5(3B) + DSTa model outperforms on specific tasks like Hellaswag, it falls short in others, including RTE and WSC. Conversely, T5(3B) + InsTa consistently enhances performance across the majority of tasks relative to T0-3B, and achieves an impressive average performance enhancement of 5.08% over T5(3B) + DSTa. This outcome suggests that instances within data samples might obstruct the extraction of representative task features, diminishing the task selector’s ability to identify related tasks effectively. Please refer to Appendix I for more experimental details.
@@ 6.2 Impact of Instruction Refinement
As mentioned in Section 4.2, our approach operates based on the premise that the instruction accurately describes the characteristics of the task, necessitating the process of instruction refinement. Table 6 demonstrates the impact of such instruction refinement on performance. The term "Unfiltered" denotes the training conducted without instruction refinement, while "Filtered" indicates the use of refined models.
Unfiltered models, while showing performance improvements in most tasks compared to T0-3B, encounter performance degradation in certain tasks, notably StoryCloze, due to the selection of irrelevant tasks. Conversely, models that use filtered and refined instructions accurately select related tasks and generally demonstrate improved performance over unfiltered models. This result emphasizes the significance of instruction quality in task selection. Furthermore, the notable performance enhancement seen in T5(3B) + InsTa underscores the substantial role of instruction quality, particularly when additional training for the alignment is applied.
@@ 6.3 Scaling Relevant Tasks
We explore how the model performance is affected by the number of selected tasks during instruction tuning. Figure 2 illustrates the average performance of T5(3B) + InsTa and T5(3B) + InsTa, as the number of selected tasks () increases. The performance of the top-35 tasks for P3 and top-756 tasks for NIV2 corresponds to the fully instruction-tuned models, representing the scores of T0-3B and Tk-Instruct-3B, respectively. On average, we observe a progressive improvement in performance up to five tasks for P3 and seventy tasks for NIV2, after which it declines, affirming the adverse impact of non-relevant tasks during training. For more analysis and detailed results, please refer to Appendix J.
@@ 7 Conclusion
In this study, we discover that selecting informative tasks for instruction tuning can be effectively achieved by exclusively using task instructions. Our experiments reveal that this method, particularly when aligned with the meta-dataset’s instructions, surpasses traditional methods which depend on data samples to determine task relevance. Beyond its robust performance, our approach’s most significant advantage is its simplicity: as long as the instruction (task description) of the target task can be described, our method can be applied. This approach marks a significant shift away from traditional methods that require exhaustive pairwise comparisons or the labor-intensive creation of data samples for new tasks. By adopting an instruction-only strategy, our method simplifies and automates the task selection process for instruction tuning, providing a more efficient and practical approach to developing models in real-world scenarios.
@@
While highlighting the effective task selection method in instruction tuning, we do not perform experimental results over the different sizes of model parameters other than the T5(3B) parameter model due to computational cost. For example, a language model bigger than 11B parameters may be less susceptible to negative transfer due to model capacity, or the effectiveness of our task selection method might stand out even more. Moreover, we only use the encoder-decoder architecture model in the paper. We leave the investigation on the decoder model, such as LLaMA 7B/13B, to our future work.
@@
Acknowledgements
We would like to express sincere thanks to Dongkyu Lee and Joel Jang for their thoughtful feedback on the paper.
@@ A.1 Task selection in Intermediate-Task Transfer Learning
Within the scope of intermediate-task transfer learning, the goal of task selection is to enhance the performance by further training the model on related intermediate tasks before fine-tuning it on a target task Phang et al. (2018); Vu et al. (2020b); Kung et al. (2021); Poth et al. (2021b). Therefore, it presupposes the availability of a labeled dataset for the target task.
Vu et al. (2020b) use BERT as a feature extractor to identify related tasks by comparing embeddings with an auxiliary task pool. In contrast, Kung et al. (2021) develop a more efficient method, training a task discriminator with just 500 samples from each dataset, thereby reducing data needs. However, creating 500 labeled samples for a target task can still be burdensome, especially in an instruction-tuning setting. Moreover, in line with Vu et al. (2020b), this approach entails a high computational cost for training task discriminator and inference on all auxiliary instances.
Poth et al. (2021b) opt for a different approach, directly measuring pairwise transfer for all datasets instead of using samples’ embedding. To reduce computational costs, they employ adapter structures. Nonetheless, this method still incurs significant costs in terms of target data construction and computational overhead because it requires training and inference across all models.
@@ A.2 Task Selection in Instruction Tuning
In task selection in instruction tuning, the objective is to select and train related tasks for improving the zero-shot performance of the target task. Unlike intermediate-task transfer learning, where samples consist solely of instances, data samples in instruction tuning include both instructions and instances. A model trained with such instructions can tackle new tasks when given new instructions, leveraging previously learned instructions as a basis.
@@
Ivison et al. (2023) propose a sample-based task selection methodology in instruction tuning. They encode all instances from a vast multitask data pool using a T5-3B model, then build a search index from the resulting representations. A key advantage of their approach is that there’s no need to train an embedding generator with the target task’s data samples, unlike task selection in intermediate-task transfer learning. However, the method still requires 1,000 unlabeled data samples from the target task. Jang et al. (2023) adopt a method similar to Ivison et al. (2023) to retrieve experts related to the target task for instruction tuning by encoding 32 data samples for each expert to calculate similarity.
@@
Zhou et al. (2023); Kim et al. (2023a) apply a methodology similar to Poth et al. (2021b) in the context of instruction tuning, performing pairwise transfer across a variety of tasks. Notably, Zhou et al. (2023) conduct pairwise transfer for all tasks in the P3 dataset, and these results are utilized as a baseline in their paper.
A.3 Baseline for Task Selection in Instruction Tuning
@@
As noted by Ivison et al. (2023), task selection in intermediate-task transfer learning has several disadvantages compared to the task selection methodologies used in instruction tuning settings. These methodologies are primarily applied to classification tasks, require a large number of labeled samples for the target task, and involve high computational costs for training a model to generate task embeddings. For these reasons, we exclude these studies in our baselines. Additionally, we used Jang et al. (2023)’s methodology as the baseline for the sample-based task selection instead of Ivison et al. (2023) because Jang et al. (2023)’s approach requires fewer data samples, making implementation easier. The characteristics and practical applicability of each methodology are summarized in the Table 7.
| Model | Time Complexity | P3 Avg. |
| T5(3B) + PE w/RoE | 53.26 | |
| T5(3B) + InsTa | 55.71 | |
| Model | Selection Time (Training) | P3 Avg. |
| T5(3B) + Pairwise Transfer | 35 * 32h | 57.86 |
| T5(3B) + InsTa | 5m | 57.97 |
Appendix B Efficiency Analysis
To evaluate the efficiency of our task selection method relative to other alternatives, we quantify it based on time and complexity. Specifically, we conduct a comparative analysis of the sample-based approach and the pairwise transfer approach using the same configurations and models as outlined in Table 2.
Sample-Based Selection vs. InsTa
Table 8 demonstrates the efficiency differences between the sample-based approach and our instruction-based approach. For this comparison, we utilize the sample-based T5(3B) + PE w/RoE and our instruction-based T5(3B) + InsTa. Both approaches employ SentenceBERT-based cosine similarity, enabling us to evaluate their efficiency by examining differences in time complexity. This time complexity incorporates both the sentence encoding process and the similarity computation. Let represent the number of training tasks, the number of evaluation tasks, the average number of instructions, and the number of data samples. The time complexity for encoding in the sample-based approach can be expressed as . In contrast, the instruction-based methodology, which does not require processing multiple data samples per instruction, has an encoding time complexity of . The time complexity for similarity computation is also derived similarly. Considering all combinations of training and evaluation tasks, the complexity for the sample-based scenario is , whereas for the instruction-based scenario, it is . This indicates that our approach is at least 32 times faster than the sample-based T5(3B) + PE w/RoE, which requires 32 samples .
Pairwise Transfer vs. InsTa
Table 9 presents the training time required for our method compared to the pairwise transfer method, which trains and evaluates every task pair. In this comparison, we analyze T5(3B) + Pairwise Transfer and T5(3B) + InsTa. Following the experimental setting on Zhou et al. (2023), with a batch size of 512 and 1000 steps, it takes approximately 32 hours on 1 A100 GPU for one task training. The P3 has 35 training datasets, and the total time required is approximately 35 * 32 hours. Conversely, our methodology takes about 5 minutes to train the SentenceBERT model on a single A100 GPU. This is only considering the training time, and the time difference becomes even greater when including inference time.
Appendix C Addressing Potential Concerns
Potential for Bias
Our task selection process, which relies solely on instructions, might initially raise concerns about potential bias. If the training data selected is biased, the tasks chosen could perpetuate this bias in the fine-tuned LLM. However, it is crucial to note that our methodology robustly counters such biases for two key reasons. Firstly, many of the publicly available instruction tuning datasets are composed of diverse datasets, which are predominantly high-quality and unbiased. Secondly, even if some of the selected datasets exhibit bias, this issue is effectively mitigated by the majority of other selected unbiased datasets. For instance, in the NIV2 experiment, we used 70 out of 756 datasets for training. Even if some datasets contained biases, the impact is neutralized by training with the remaining datasets. Our experiments support this approach, as we confirmed no performance degradation due to bias across a total of 791 training datasets and 85 evaluation tasks (P3: 11, NIV2: 33, BB: 14, BBH: 27). These results demonstrate that our approach ensures robustness against potential biases.
Limited Generalizability
Our task selection method relies on the quality and completeness of instructions for the target task. Therefore, there may be concerns that if the instructions are inaccurate, performance could deteriorate. However, it’s important to note that our methodology can effectively operate with just one well-written instruction. Creating such an instruction is not typically burdensome. Moreover, should challenges arise in defining the instruction, they can be readily resolved using models like GPT-4 (please refer to Table 13 for the GPT-4 query). For instance, for the BBH datasets, which initially lacked instructions, we utilized GPT-4 to generate one instruction per task for evaluation purposes. The results demonstrated significant performance improvements in most tasks (see Table 4). Overall, while our methodology cannot completely eliminate issues with instruction quality, these limitations do not severely restrict its generalizability and can be easily alleviated.
Appendix D Off-the-shelf Embedding Model
@@
During training, we employ a learning rate of 1e-6 for the P3 dataset and 1e-5 for the NIV2 dataset. The remaining training configurations align with those described in Reimers and Gurevych (2019). Specifically, we conduct training for five epochs and select the checkpoint corresponding to the best validation performance for subsequent analysis and evaluation.
Appendix E Training Details
Following Sanh et al. (2022), we conduct our experiments in a held-out setting. Specifically, we exclude the task clusters of the evaluation tasks from the training dataset pool and then train our model, subsequently evaluating it on the excluded held-out tasks. For instance, in the case of the P3 dataset, we remove task clusters such as sentence completion, NLI, coreference resolution, and word sense disambiguation from the training set and train only on the selected tasks from the remaining datasets. It ensures that the evaluation set remains entirely unseen, thereby achieving a true zero-shot setting.
Additionally, SentenceBert training for InsTa follows the same principle. We exclude all instructions from the evaluation clusters for task selector training and train using only instructions from the training dataset pool, thus preventing data contamination issues.
@@ Appendix E Training Details
Appendix F Examples and Formulation of Instructions of P3 and NIV2.
F.1 Examples of P3 and NIV2 Instructions.
@@ F.3 NIV2 Input Formats
Appendix G GPT-4 API Generation
G.1 NIV2 Instruction Generation
G.2 Big-Bench Hard Instruction Generation
The BBH tasks don’t have refined Task Definition like NIV2 datasets. In order to select relevant tasks using NIV2 datasets, and to compare the performance with Tk-Instruct, we generate instruction for BBH tasks using GPT-4. Table 15 represents API query used for instruction generation, and Table 16, 17 and 18 show generated 27 instructions for each task.
Appendix H Pairwise Transfer Task Selection vs. Instruction-Based Task Selection
@@ G.2 Big-Bench Hard Instruction Generation
Appendix I Experimental Detail for Data Sample-based Task Selection
@@
Appendix J Analysis on Selecting Top-k Relevant Tasks
@@
Appendix K Scaling The Number of Instructions in Alignment
@@
Appendix L Learning Different Formats of Meta Datasets and Its Impact on Performance